# Detection of Gravitational Anomaly at Low Acceleration from a Highest-quality Sample of 36 Wide Binaries with Accurate 3D Velocities

### v1

@@ E.5 Speckle Holography
Gaia DR3 3270823243234194432 2024.8618 1.30 0.93
Gaia DR3 3270823243234194944 2024.8618 1.09 0.73
Gaia DR3 3279804084274429440 2024.8620 2.52 2.10
Gaia DR3 3279804844484336128 2024.8620 0.61 0.36
Gaia DR3 3285744612456150016 2024.8620 0.32 0.23
Gaia DR3 3285755951169814528 2024.8620 0.42 0.28
Gaia DR3 3288572968680438528 2024.8621 0.22 0.15
Gaia DR3 3288572968680438912 2024.8621 0.20 0.14
Gaia DR3 3290322734060874752 2024.8649 0.63 0.37
Gaia DR3 3290322738355773696 2024.8649 0.81 0.42
@@
Gaia DR3 3290540746599303296 2024.8649 0.63 0.37
Gaia DR3 3308694065827291648 2024.8621 0.47 0.31
Gaia DR3 3308694065827292160 2024.8621 0.71 0.39
Gaia DR3 3338164860103556352 2024.8649 0.40 0.28
Gaia DR3 3338164860104174592 2024.8649 1.23 0.73
Gaia DR3 3344333085975753216 2024.8623 0.63 0.37
Gaia DR3 3344333292134182400 2024.8623 1.06 0.62
Gaia DR3 3346799805954385408 2024.8649 0.48 0.31
Gaia DR3 3346799805954387456 2024.8649 0.19 0.14
Gaia DR3 3360175090027418496 2024.8624 0.44 0.29
Gaia DR3 3360175090027418880 2024.8624 1.80 1.48
Gaia DR3 3365948002813451648 2024.8623 1.40 1.00
Gaia DR3 3365948002813451776 2024.8623 1.45 1.08
Gaia DR3 3371529368651025664 2024.1278 0.40 0.28
Gaia DR3 3371529467432172672 2024.1279 0.48 0.33
Gaia DR3 3399607116051404032 2024.8649 0.59 0.36
@@
Gaia DR3 3400637667683964544 2024.8649 0.55 0.34
Gaia DR3 3400637702043702656 2024.8649 0.41 0.27
This is a triple system. Gaia DR3 3402090259984528768 2024.1278 0.40 0.28
@@
This is a triple system. Gaia DR3 3402090466142958464 2024.1278 0.51 0.33
@@
Gaia DR3 3410692942038997120 2024.8621 0.39 0.27
Gaia DR3 3410692942038997760 2024.8621 0.43 0.29
Gaia DR3 3417608762800814080 2024.8650 0.32 0.23
Gaia DR3 3417608767098501120 2024.8650 0.37 0.26
Gaia DR3 3418915433586276096 2024.8649 0.33 0.24
Gaia DR3 3418915433588729856 2024.8649 1.06 0.56
Gaia DR3 3451266742171824640 2024.8650 0.20 0.15
Gaia DR3 3451267120128948992 2024.1278 0.19 0.15
Gaia DR3 3593154766362734336 2025.2179 0.31 0.23
Gaia DR3 3593154766362734464 2025.2179 0.31 0.23
Gaia DR3 3609320267350909696 2025.4392 0.47 0.32
Gaia DR3 3609320503573212416 2025.4392 0.53 0.36
Gaia DR3 3609413309226560768 2025.4392 1.15 0.85
Gaia DR3 3609413485320886528 2025.4392 0.35 0.26
Gaia DR3 3618102204160590080 2025.4446 0.45 0.32
Gaia DR3 3618102204160590208 2025.2181 0.27 0.20
Gaia DR3 3627968323859543680 2025.4446 0.34 0.26
Gaia DR3 3627968328155313024 2025.4446 1.40 1.06
Gaia DR3 3629449713915122688 2025.4446 0.25 0.19
Gaia DR3 3629449713915123072 2025.4446 1.00 0.65
Gaia DR3 3636691509812473856 2025.4392 0.26 0.18
Gaia DR3 3636714977513779712 2025.4392 0.31 0.22
Gaia DR3 3639520621950395776 2025.4447 0.12 0.10
Gaia DR3 3639520621950395904 2025.4447 0.12 0.10
The third component has no Gaia identification; Gaia DR3 3641315024925921024 2025.4393 0.27 0.20
@@
The third component has no Gaia identification; Gaia DR3 3641315024927514880 2025.4393 0.69 0.40
@@
Gaia DR3 3666150003300751744 2025.4448 0.22 0.16
Gaia DR3 3666150003300991744 2025.4448 0.25 0.18
Gaia DR3 367020421620737536 2024.6815 0.28 0.20
Gaia DR3 367020421621893120 2024.6815 0.26 0.19
Gaia DR3 3677417214346066048 2025.2180 0.34 0.25
Gaia DR3 3677417214346902144 2025.2180 0.58 0.37
Gaia DR3 3680089852236243584 2025.2180 0.23 0.17
Gaia DR3 3680089852236243712 2025.2180 0.25 0.19
Gaia DR3 3686259727375524864 2025.4392 0.43 0.29
Gaia DR3 3686261307923490688 2025.4392 0.25 0.18
Gaia DR3 3717756234385505792 2025.4446 0.19 0.14
Gaia DR3 3717756337464721152 2025.4364 0.76 0.50
Gaia DR3 3718581417862169344 2025.4364 0.35 0.26
Gaia DR3 3718581520941384320 2025.4447 0.41 0.29
Gaia DR3 372066046121799168 2024.6870 0.00 0.00
Gaia DR3 372066076185339648 2024.6869 0.26 0.18
Gaia DR3 3724521696934640384 2025.4447 0.20 0.14
Gaia DR3 3724521701229940096 2025.4447 0.34 0.24
Gaia DR3 3738783878870503424 2025.4365 0.34 0.25
Gaia DR3 3738783878870503552 2025.4365 0.77 0.53
Gaia DR3 3739054324371407232 2025.4447 0.43 0.31
Gaia DR3 3739054324371407488 2025.4447 0.24 0.18
Gaia DR3 374529119673672576 2024.6870 0.38 0.25
Gaia DR3 374529123966850944 2024.6870 0.29 0.20
Gaia DR3 3768566346037170304 2025.2178 0.34 0.25
Gaia DR3 3768566350332212352 2025.2178 0.34 0.25
Gaia DR3 377244848968006016 2024.6816 0.26 0.18
Gaia DR3 377245055126435072 2024.6816 0.22 0.15
Gaia DR3 3792739899447945216 2024.1285 0.73 0.52
Gaia DR3 3792740006822516608 2024.1285 1.10 0.89
Gaia DR3 3793106419072038272 2025.2178 0.26 0.19
Gaia DR3 3793107930900527616 2025.2178 0.20 0.14
Gaia DR3 3838561363635304448 2024.8679 0.35 0.25
Gaia DR3 3838561363635829632 2024.8679 0.34 0.24
Gaia DR3 3841514200895729152 2024.8679 0.37 0.24
Gaia DR3 3841514200898362368 2024.8679 0.33 0.23
Gaia DR3 3890860179670959104 2025.2178 0.25 0.18
Gaia DR3 3890860183966486656 2025.2178 0.24 0.17
The third component is DR3 3895404602964088448. Gaia DR3 3895404602962524544 2025.2179 0.20 0.15
@@
The third component is DR3 3895404602964088448. Gaia DR3 3895404602964117504 2025.2179 0.32 0.23
@@
Gaia DR3 3930274446007361664 2025.4365 0.21 0.16
Gaia DR3 3930274613510165504 2025.4365 0.17 0.13
This is a multi-component system. Gaia DR3 3936935283153115008 2025.4419 0.24 0.16
@@
This is a multi-component system. Gaia DR3 3936947278997113216 2025.4419 1.03 0.81
@@
Gaia DR3 3945118265299248128 2025.4447 0.14 0.11
Gaia DR3 3945118643256370688 2025.4447 0.16 0.12
Gaia DR3 3974007108685133056 2024.1285 1.00 0.70
Gaia DR3 3974007585425593472 2025.2179 0.68 0.40
Gaia DR3 3975566491051012480 2025.2179 0.25 0.18
Gaia DR3 3975566491051012608 2025.2179 0.25 0.18
Gaia DR3 4007902062872083456 2024.1285 0.41 0.30
Gaia DR3 4007902269030513920 2024.1285 0.95 0.70
Gaia DR3 4025108384759142272 2025.2179 0.60 0.37
Gaia DR3 405028575795255808 2024.6870 0.26 0.19
Gaia DR3 405028683170219392 2024.6870 0.31 0.22
Gaia DR3 4103210548364197376 2025.4452 0.23 0.17
Gaia DR3 4103210582723957120 2025.4452 0.31 0.24
Gaia DR3 418549201560407040 2024.6815 0.45 0.32
Gaia DR3 418549205864071296 2024.6815 0.43 0.30
Gaia DR3 4197023904013455872 2025.4371 0.19 0.13
Gaia DR3 4197023904014523776 2025.4371 0.16 0.12
Gaia DR3 4218533748765026560 2025.4373 0.32 0.23
Gaia DR3 4218533959216594176 2025.4373 0.77 0.44
Gaia DR3 4230699329529382400 2025.4373 0.28 0.20
Gaia DR3 4230699363889120128 2025.4373 0.18 0.13
Gaia DR3 4231853855392895104 2024.6783 0.94 0.48
Gaia DR3 4231853855392895360 2024.6783 1.24 0.70
Gaia DR3 4232150409999693056 2024.6783 0.42 0.29
Gaia DR3 4232150410001183872 2024.6783 0.42 0.29
Gaia DR3 4243504314963260672 2024.6780 0.36 0.25
Gaia DR3 4243504314963261696 2024.6780 1.10 0.68
Gaia DR3 4245247040891396480 2024.6782 1.53 1.23
Gaia DR3 4245247247049828480 2024.6782 0.41 0.29
Gaia DR3 4249024035142729728 2024.6782 1.19 0.74
Gaia DR3 4249024035142730368 2024.6782 0.38 0.26
Gaia DR3 428811394564909568 2024.6814 0.24 0.17
Gaia DR3 428811428924644352 2024.6814 0.32 0.23
Gaia DR3 4292414749731991808 2025.4371 0.31 0.21
Gaia DR3 4292415883603395456 2025.4371 0.46 0.30
Gaia DR3 4335277282059878656 2025.4343 0.53 0.36
Gaia DR3 4335277282066850048 2025.4343 0.56 0.37
Gaia DR3 4350763323519241600 2025.4343 0.63 0.39
Gaia DR3 4350763323519241984 2025.4343 0.80 0.45
Gaia DR3 4353764371788098944 2025.4369 0.36 0.26
Gaia DR3 4353764436210206080 2025.4369 0.26 0.19
Gaia DR3 4368065856969737728 2025.4396 0.20 0.15
Gaia DR3 4368065921391001600 2025.4396 0.77 0.42
Gaia DR3 4385319393432940544 2025.4395 0.81 0.44
Gaia DR3 4385319427793947648 2025.4395 0.32 0.22
Gaia DR3 4390274857979921408 2025.4396 0.47 0.31
Gaia DR3 4390274857979922432 2025.4396 0.26 0.19
Gaia DR3 4395522998779760128 2025.4367 0.29 0.21
Gaia DR3 4395523033138822656 2025.4367 0.29 0.21
Gaia DR3 4404091664690806016 2025.4340 0.23 0.17
Gaia DR3 4404091664691143168 2025.4340 0.28 0.20
Gaia DR3 4407541313703639424 2025.4369 1.07 0.69
Gaia DR3 4407544268641139712 2025.4369 1.12 0.74
Gaia DR3 441620979895171840 2024.8619 0.30 0.22
Gaia DR3 441620979901788544 2024.8619 0.64 0.39
Gaia DR3 4430185034123000960 2025.4341 0.26 0.19
Gaia DR3 4430185068482324864 2025.4367 0.24 0.17
Gaia DR3 4435683451255623808 2025.4395 0.24 0.17
Gaia DR3 4435689739087756800 2025.4395 0.11 0.10
Gaia DR3 4441920942064076288 2025.4395 0.47 0.31
Gaia DR3 4441920946360132096 2025.4395 0.37 0.25
Gaia DR3 4445849153513150080 2025.4450 0.25 0.18
Gaia DR3 4445849157808061312 2025.4450 0.25 0.18
Gaia DR3 4451719690908056192 2025.4341 0.28 0.20
Gaia DR3 4451731819895701376 2025.4341 1.02 0.64
Gaia DR3 4454413455738305792 2025.4341 0.60 0.38
Gaia DR3 4454413460037125888 2025.4341 0.62 0.38
Gaia DR3 4466613022581649152 2025.4343 0.78 0.48
Gaia DR3 4466613022581649664 2025.4343 0.64 0.41
Gaia DR3 4471562916556377856 2025.4397 0.27 0.20
Gaia DR3 4471658131695403264 2025.4397 0.77 0.41
Gaia DR3 4474801218817508864 2025.4370 0.72 0.41
Gaia DR3 4474801257476653952 2025.4370 0.34 0.24
The third component is DR3 4493544082437142784 , Gaia DR3 4493544082437142400 2025.4370 0.84 0.49
@@
The third component is DR3 4493544082437142784 , Gaia DR3 4493544185516358656 2025.4370 0.84 0.47
@@
Gaia DR3 4502416488439872128 2025.4370 0.24 0.17
Gaia DR3 4502416488439872256 2025.4370 0.26 0.18
Gaia DR3 4513425898394508288 2025.4398 0.31 0.22
Gaia DR3 4513425932752582144 2025.4398 0.42 0.28
Gaia DR3 4513771449965175296 2025.4398 0.60 0.36
Gaia DR3 4513771832248074240 2025.4398 1.31 0.92
Gaia DR3 4541750383049979904 2025.4396 0.38 0.26
Gaia DR3 4541753303627759104 2025.4396 0.24 0.17
Gaia DR3 4544016820111855360 2025.4370 0.20 0.14
Gaia DR3 4544017129349496832 2025.4370 0.32 0.22
Gaia DR3 4558397710654524288 2025.4344 0.26 0.20
Gaia DR3 4558397710655578368 2025.4344 0.38 0.28
Gaia DR3 4571298692894669824 2025.4344 0.77 0.44
Gaia DR3 4571298692894670336 2025.4344 0.91 0.50
Gaia DR3 4572807566444752128 2025.4395 0.32 0.22
Gaia DR3 4577271961615916032 2025.4370 0.33 0.23
Gaia DR3 4577272270853565952 2025.4370 0.21 0.15
Gaia DR3 4584470704757139968 2025.4397 0.46 0.30
Gaia DR3 4584470739116876800 2025.4397 0.35 0.23
Gaia DR3 4589844945796230016 2025.4452 0.17 0.12
Gaia DR3 4589850065397245824 2025.4452 0.30 0.21
Gaia DR3 459661938487951360 2024.8617 0.49 0.32
Gaia DR3 459661938487952000 2024.8617 0.98 0.56
Gaia DR3 4599117333509619200 2025.4397 1.12 0.63
Gaia DR3 4599118119486450688 2025.4397 0.70 0.37
Gaia DR3 4599984504586131456 2025.4397 0.23 0.17
Gaia DR3 4599984642025088128 2025.4397 0.14 0.11
Gaia DR3 4600068239268380928 2025.4370 0.00 0.00
Gaia DR3 4600068239268381312 2025.4370 0.27 0.19
Gaia DR3 460621193663868160 2024.8617 0.34 0.25
Gaia DR3 460621296743080960 2024.8617 1.24 0.84
Gaia DR3 4724021644644224 2024.8617 1.23 0.86
Gaia DR3 4725155516009856 2024.8617 1.15 0.77
Gaia DR3 5110662129034491904 2024.8619 1.02 0.62
Gaia DR3 5110662129034492032 2024.8619 0.61 0.37
Gaia DR3 5114544745110388736 2024.8619 0.70 0.41
Gaia DR3 5114547700047887360 2024.8619 0.88 0.50
Gaia DR3 5166931251492281472 2024.8619 1.24 0.89

### v2

@@ E.5 Speckle Holography
Gaia DR3 3270823243234194432 2024.8618 0.650 0.465
Gaia DR3 3270823243234194944 2024.8618 0.545 0.365
Gaia DR3 3279804084274429440 2024.862 1.260 1.050
Gaia DR3 3279804844484336128 2024.862 0.305 0.180
Gaia DR3 3285744612456150016 2024.862 0.160 0.115
Gaia DR3 3285755951169814528 2024.862 0.210 0.140
Gaia DR3 3288572968680438528 2024.8621 0.110 0.075
Gaia DR3 3288572968680438912 2024.8621 0.100 0.070
Gaia DR3 3290322734060874752 2024.8649 0.315 0.185
Gaia DR3 3290322738355773696 2024.8649 0.405 0.210
@@
Gaia DR3 3290540746599303296 2024.8649 0.315 0.185
Gaia DR3 3308694065827291648 2024.8621 0.235 0.155
Gaia DR3 3308694065827292160 2024.8621 0.355 0.195
Gaia DR3 3338164860103556352 2024.8649 0.200 0.140
Gaia DR3 3338164860104174592 2024.8649 0.615 0.365
Gaia DR3 3344333085975753216 2024.8623 0.315 0.185
Gaia DR3 3344333292134182400 2024.8623 0.530 0.310
Gaia DR3 3346799805954385408 2024.8649 0.240 0.155
Gaia DR3 3346799805954387456 2024.8649 0.095 0.070
Gaia DR3 3360175090027418496 2024.8624 0.220 0.145
Gaia DR3 3360175090027418880 2024.8624 0.900 0.740
Gaia DR3 3365948002813451648 2024.8623 0.700 0.500
Gaia DR3 3365948002813451776 2024.8623 0.725 0.540
Gaia DR3 3371529368651025664 2024.1278 0.200 0.140
Gaia DR3 3371529467432172672 2024.1279 0.240 0.165
Gaia DR3 3399607116051404032 2024.8649 0.295 0.180
@@
Gaia DR3 3400637667683964544 2024.8649 0.275 0.170
Gaia DR3 3400637702043702656 2024.8649 0.205 0.135
This is a triple system. Gaia DR3 3402090259984528768 2024.1278 0.200 0.140
@@
This is a triple system. Gaia DR3 3402090466142958464 2024.1278 0.255 0.165
@@
Gaia DR3 3410692942038997120 2024.8621 0.195 0.135
Gaia DR3 3410692942038997760 2024.8621 0.215 0.145
Gaia DR3 3417608762800814080 2024.865 0.160 0.115
Gaia DR3 3417608767098501120 2024.865 0.185 0.130
Gaia DR3 3418915433586276096 2024.8649 0.165 0.120
Gaia DR3 3418915433588729856 2024.8649 0.530 0.280
Gaia DR3 3451266742171824640 2024.865 0.100 0.075
Gaia DR3 3451267120128948992 2024.1278 0.095 0.075
Gaia DR3 3593154766362734336 2025.2179 0.155 0.115
Gaia DR3 3593154766362734464 2025.2179 0.155 0.115
Gaia DR3 3609320267350909696 2025.4392 0.235 0.160
Gaia DR3 3609320503573212416 2025.4392 0.265 0.180
Gaia DR3 3609413309226560768 2025.4392 0.575 0.425
Gaia DR3 3609413485320886528 2025.4392 0.175 0.130
Gaia DR3 3618102204160590080 2025.4446 0.225 0.160
Gaia DR3 3618102204160590208 2025.2181 0.135 0.100
Gaia DR3 3627968323859543680 2025.4446 0.170 0.130
Gaia DR3 3627968328155313024 2025.4446 0.700 0.530
Gaia DR3 3629449713915122688 2025.4446 0.125 0.095
Gaia DR3 3629449713915123072 2025.4446 0.500 0.325
Gaia DR3 3636691509812473856 2025.4392 0.130 0.090
Gaia DR3 3636714977513779712 2025.4392 0.155 0.110
Gaia DR3 3639520621950395776 2025.4447 0.060 0.050
Gaia DR3 3639520621950395904 2025.4447 0.060 0.050
The third component has no Gaia identification; Gaia DR3 3641315024925921024 2025.4393 0.135 0.100
@@
The third component has no Gaia identification; Gaia DR3 3641315024927514880 2025.4393 0.345 0.200
@@
Gaia DR3 3666150003300751744 2025.4448 0.110 0.080
Gaia DR3 3666150003300991744 2025.4448 0.125 0.090
Gaia DR3 367020421620737536 2024.6815 0.140 0.100
Gaia DR3 367020421621893120 2024.6815 0.130 0.095
Gaia DR3 3677417214346066048 2025.218 0.170 0.125
Gaia DR3 3677417214346902144 2025.218 0.290 0.185
Gaia DR3 3680089852236243584 2025.218 0.115 0.085
Gaia DR3 3680089852236243712 2025.218 0.125 0.095
Gaia DR3 3686259727375524864 2025.4392 0.215 0.145
Gaia DR3 3686261307923490688 2025.4392 0.125 0.090
Gaia DR3 3717756234385505792 2025.4446 0.095 0.070
Gaia DR3 3717756337464721152 2025.4364 0.380 0.250
Gaia DR3 3718581417862169344 2025.4364 0.175 0.130
Gaia DR3 3718581520941384320 2025.4447 0.205 0.145
Too weak Gaia DR3 372066046121799168 2024.687 >2.>2. >2.>2.
Gaia DR3 372066076185339648 2024.6869 0.130 0.090
Gaia DR3 3724521696934640384 2025.4447 0.100 0.070
Gaia DR3 3724521701229940096 2025.4447 0.170 0.120
Gaia DR3 3738783878870503424 2025.4365 0.170 0.125
Gaia DR3 3738783878870503552 2025.4365 0.385 0.265
Gaia DR3 3739054324371407232 2025.4447 0.215 0.155
Gaia DR3 3739054324371407488 2025.4447 0.120 0.090
Gaia DR3 374529119673672576 2024.687 0.190 0.125
Gaia DR3 374529123966850944 2024.687 0.145 0.100
Gaia DR3 3768566346037170304 2025.2178 0.170 0.125
Gaia DR3 3768566350332212352 2025.2178 0.170 0.125
Gaia DR3 377244848968006016 2024.6816 0.130 0.090
Gaia DR3 377245055126435072 2024.6816 0.110 0.075
Gaia DR3 3792739899447945216 2024.1285 0.365 0.260
Gaia DR3 3792740006822516608 2024.1285 0.550 0.445
Gaia DR3 3793106419072038272 2025.2178 0.130 0.095
Gaia DR3 3793107930900527616 2025.2178 0.100 0.070
Gaia DR3 3838561363635304448 2024.8679 0.175 0.125
Gaia DR3 3838561363635829632 2024.8679 0.170 0.120
Gaia DR3 3841514200895729152 2024.8679 0.185 0.120
Gaia DR3 3841514200898362368 2024.8679 0.165 0.115
Gaia DR3 3890860179670959104 2025.2178 0.125 0.090
Gaia DR3 3890860183966486656 2025.2178 0.120 0.085
The third component is DR3 3895404602964088448. Gaia DR3 3895404602962524544 2025.2179 0.100 0.075
@@
The third component is DR3 3895404602964088448. Gaia DR3 3895404602964117504 2025.2179 0.160 0.115
@@
Gaia DR3 3930274446007361664 2025.4365 0.105 0.080
Gaia DR3 3930274613510165504 2025.4365 0.085 0.065
This is a multi-component system. Gaia DR3 3936935283153115008 2025.4419 0.120 0.080
@@
This is a multi-component system. Gaia DR3 3936947278997113216 2025.4419 0.515 0.405
@@
Gaia DR3 3945118265299248128 2025.4447 0.070 0.055
Gaia DR3 3945118643256370688 2025.4447 0.080 0.060
Gaia DR3 3974007108685133056 2024.1285 0.500 0.350
Gaia DR3 3974007585425593472 2025.2179 0.340 0.200
Gaia DR3 3975566491051012480 2025.2179 0.125 0.090
Gaia DR3 3975566491051012608 2025.2179 0.125 0.090
Gaia DR3 4007902062872083456 2024.1285 0.205 0.150
Gaia DR3 4007902269030513920 2024.1285 0.475 0.350
Gaia DR3 4025108384759142272 2025.2179 0.300 0.185
Gaia DR3 405028575795255808 2024.687 0.130 0.095
Gaia DR3 405028683170219392 2024.687 0.155 0.110
Gaia DR3 4103210548364197376 2025.4452 0.115 0.085
Gaia DR3 4103210582723957120 2025.4452 0.155 0.120
Gaia DR3 418549201560407040 2024.6815 0.225 0.160
Gaia DR3 418549205864071296 2024.6815 0.215 0.150
Gaia DR3 4197023904013455872 2025.4371 0.095 0.065
Gaia DR3 4197023904014523776 2025.4371 0.080 0.060
Gaia DR3 4218533748765026560 2025.4373 0.160 0.115
Gaia DR3 4218533959216594176 2025.4373 0.385 0.220
Gaia DR3 4230699329529382400 2025.4373 0.140 0.100
Gaia DR3 4230699363889120128 2025.4373 0.090 0.065
Gaia DR3 4231853855392895104 2024.6783 0.470 0.240
Gaia DR3 4231853855392895360 2024.6783 0.620 0.350
Gaia DR3 4232150409999693056 2024.6783 0.210 0.145
Gaia DR3 4232150410001183872 2024.6783 0.210 0.145
Gaia DR3 4243504314963260672 2024.678 0.180 0.125
Gaia DR3 4243504314963261696 2024.678 0.550 0.340
Gaia DR3 4245247040891396480 2024.6782 0.765 0.615
Gaia DR3 4245247247049828480 2024.6782 0.205 0.145
Gaia DR3 4249024035142729728 2024.6782 0.595 0.370
Gaia DR3 4249024035142730368 2024.6782 0.190 0.130
Gaia DR3 428811394564909568 2024.6814 0.120 0.085
Gaia DR3 428811428924644352 2024.6814 0.160 0.115
Gaia DR3 4292414749731991808 2025.4371 0.155 0.105
Gaia DR3 4292415883603395456 2025.4371 0.230 0.150
Gaia DR3 4335277282059878656 2025.4343 0.265 0.180
Gaia DR3 4335277282066850048 2025.4343 0.280 0.185
Gaia DR3 4350763323519241600 2025.4343 0.315 0.195
Gaia DR3 4350763323519241984 2025.4343 0.400 0.225
Gaia DR3 4353764371788098944 2025.4369 0.180 0.130
Gaia DR3 4353764436210206080 2025.4369 0.130 0.095
Gaia DR3 4368065856969737728 2025.4396 0.100 0.075
Gaia DR3 4368065921391001600 2025.4396 0.385 0.210
Gaia DR3 4385319393432940544 2025.4395 0.405 0.220
Gaia DR3 4385319427793947648 2025.4395 0.160 0.110
Gaia DR3 4390274857979921408 2025.4396 0.235 0.155
Gaia DR3 4390274857979922432 2025.4396 0.130 0.095
Gaia DR3 4395522998779760128 2025.4367 0.145 0.105
Gaia DR3 4395523033138822656 2025.4367 0.145 0.105
Gaia DR3 4404091664690806016 2025.434 0.115 0.085
Gaia DR3 4404091664691143168 2025.434 0.140 0.100
Gaia DR3 4407541313703639424 2025.4369 0.535 0.345
Gaia DR3 4407544268641139712 2025.4369 0.560 0.370
Gaia DR3 441620979895171840 2024.8619 0.150 0.110
Gaia DR3 441620979901788544 2024.8619 0.320 0.195
Gaia DR3 4430185034123000960 2025.4341 0.130 0.095
Gaia DR3 4430185068482324864 2025.4367 0.120 0.085
Gaia DR3 4435683451255623808 2025.4395 0.120 0.085
Gaia DR3 4435689739087756800 2025.4395 0.055 0.050
Gaia DR3 4441920942064076288 2025.4395 0.235 0.155
Gaia DR3 4441920946360132096 2025.4395 0.185 0.125
Gaia DR3 4445849153513150080 2025.445 0.125 0.090
Gaia DR3 4445849157808061312 2025.445 0.125 0.090
Gaia DR3 4451719690908056192 2025.4341 0.140 0.100
Gaia DR3 4451731819895701376 2025.4341 0.510 0.320
Gaia DR3 4454413455738305792 2025.4341 0.300 0.190
Gaia DR3 4454413460037125888 2025.4341 0.310 0.190
Gaia DR3 4466613022581649152 2025.4343 0.390 0.240
Gaia DR3 4466613022581649664 2025.4343 0.320 0.205
Gaia DR3 4471562916556377856 2025.4397 0.135 0.100
Gaia DR3 4471658131695403264 2025.4397 0.385 0.205
Gaia DR3 4474801218817508864 2025.437 0.360 0.205
Gaia DR3 4474801257476653952 2025.437 0.170 0.120
The third component is DR3 4493544082437142784 , Gaia DR3 4493544082437142400 2025.437 0.420 0.245
@@
The third component is DR3 4493544082437142784 , Gaia DR3 4493544185516358656 2025.437 0.420 0.235
@@
Gaia DR3 4502416488439872128 2025.437 0.120 0.085
Gaia DR3 4502416488439872256 2025.437 0.130 0.090
Gaia DR3 4513425898394508288 2025.4398 0.155 0.110
Gaia DR3 4513425932752582144 2025.4398 0.210 0.140
Gaia DR3 4513771449965175296 2025.4398 0.300 0.180
Gaia DR3 4513771832248074240 2025.4398 0.655 0.460
Gaia DR3 4541750383049979904 2025.4396 0.190 0.130
Gaia DR3 4541753303627759104 2025.4396 0.120 0.085
Gaia DR3 4544016820111855360 2025.437 0.100 0.070
Gaia DR3 4544017129349496832 2025.437 0.160 0.110
Gaia DR3 4558397710654524288 2025.4344 0.130 0.100
Gaia DR3 4558397710655578368 2025.4344 0.190 0.140
Gaia DR3 4571298692894669824 2025.4344 0.385 0.220
Gaia DR3 4571298692894670336 2025.4344 0.455 0.250
Gaia DR3 4572807566444752128 2025.4395 0.160 0.110
Gaia DR3 4577271961615916032 2025.437 0.165 0.115
Gaia DR3 4577272270853565952 2025.437 0.105 0.075
Gaia DR3 4584470704757139968 2025.4397 0.230 0.150
Gaia DR3 4584470739116876800 2025.4397 0.175 0.115
Gaia DR3 4589844945796230016 2025.4452 0.085 0.060
Gaia DR3 4589850065397245824 2025.4452 0.150 0.105
Gaia DR3 459661938487951360 2024.8617 0.245 0.160
Gaia DR3 459661938487952000 2024.8617 0.490 0.280
Gaia DR3 4599117333509619200 2025.4397 0.560 0.315
Gaia DR3 4599118119486450688 2025.4397 0.350 0.185
Gaia DR3 4599984504586131456 2025.4397 0.115 0.085
Gaia DR3 4599984642025088128 2025.4397 0.070 0.055
Too weak Gaia DR3 4600068239268380928 2025.437 >2.>2. >2.>2.
Gaia DR3 4600068239268381312 2025.437 0.135 0.095
Gaia DR3 460621193663868160 2024.8617 0.170 0.125
Gaia DR3 460621296743080960 2024.8617 0.620 0.420
Gaia DR3 4724021644644224 2024.8617 0.615 0.430
Gaia DR3 4725155516009856 2024.8617 0.575 0.385
Gaia DR3 5110662129034491904 2024.8619 0.510 0.310
Gaia DR3 5110662129034492032 2024.8619 0.305 0.185
Gaia DR3 5114544745110388736 2024.8619 0.350 0.205
Gaia DR3 5114547700047887360 2024.8619 0.440 0.250
Gaia DR3 5166931251492281472 2024.8619 0.620 0.445